The cost in dollars of manufacturing (w) wing nuts is given by the expression 0.07w + 13.3.
Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?

Found 2 solutions by fractalier, venugopalramana:
Answer by fractalier(2101)

The cost function can be written as
c(w) = .07w + 13.3
so that for 375 wing nuts, we have a cost
c(375) = (.07)(375) + 13.3 =
$39.55
The average is found by
$39.55 / 375 = $ .1055 or 10.55 cents each

Answer by venugopalramana(3286)

The cost in dollars of manufacturing (w) wing nuts is given by the expression 0.07w + 13.3.
Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?
COST OF 375 WING NUTS = 0.07*375+13.3=26.25+13.3=39.55
AVERAGE COST OF ONE WING NUT = 39.55/375=0.1054